# Fast Training of Sinusoidal Neural Fields via Scaling Initialization

### v1

@@ Abstract
Neural fields are an emerging paradigm that represent data as continuous functions parameterized by neural networks. Despite many advantages, neural fields often have a high training cost, which prevents a broader adoption. In this paper, we focus on a popular family of neural fields, called sinusoidal neural fields (SNFs), and study how it should be initialized to maximize the training speed. We find that the standard initialization scheme for SNFs—designed based on the signal propagation principle—is suboptimal. In particular, we show that by simply multiplying each weight (except for the last layer) by a constant, we can accelerate SNF training by 10$\times$. This method, coined weight scaling, consistently provides a significant speedup over various data domains, allowing the SNFs to train faster than more recently proposed architectures. To understand why the weight scaling works well, we conduct extensive theoretical and empirical analyses which reveal that the weight scaling not only resolves the spectral bias quite effectively but also enjoys a well-conditioned optimization trajectory.
@@ 1 Introduction
For neural fields, the training speed is of vital importance. Representing each datum as an NF requires tedious training of a neural network, which can take up to several hours of GPU-based training (Mildenhall et al., 2020). This computational burden becomes a critical obstacle toward adoption to tasks that involve large-scale data, such as NF-based inference or generation (Ma et al., 2024; Papa et al., 2024). To address this issue, many research efforts have been taken to accelerate NF training, including fast-trainable NF architectures (Sitzmann et al., 2020b; Müller et al., 2022), meta-learning (Sitzmann et al., 2020a; Chen & Wang, 2022), and data transformations (Seo et al., 2024).
@@
Despite such efforts, a critical aspect of NF training remains understudied: How should we initialize NFs for the fastest training? While the importance of a good initialization has been much contemplated in classic deep learning literature (Glorot & Bengio, 2010; Zhang et al., 2019), such understandings do not immediately answer our question, for many reasons. First, conventional works mostly focus on how initialization affects the signal propagation properties of deep networks with hundreds of layers, while most NFs use shallow models with only a few layers (Mildenhall et al., 2020). Second, conventional works aim to maximize the model quality at convergence, while in NFs it is often more important how fast the network achieves certain fidelity criterion (Müller et al., 2022). Lastly, NFs put great emphasis on how the model performs on seen inputs (i.e., training loss) and leaves how well it interpolates between coordinates (i.e., generalizability) as a secondary issue, while conventional neural networks care exclusively on the test performance.
@@
To fill this gap, in this paper, we investigate how the initialization affects the training efficiency of neural fields. In particular, we focus on initializing sinusoidal neural field (SNF) architectures, i.e., multilayer perceptrons with sinusoidal activation functions (Sitzmann et al., 2020b). SNF is one of the most widely used NFs, due to its versatility and parameter-efficiency; the architecture works as a strong baseline in a wide range of data modalities (Grattarola & Vandergheynst, 2022; Kim et al., 2024; Rußwurm et al., 2024) and is used as an indispensable building block of many state-of-the-art neural field frameworks (Guo et al., 2023; Schwarz et al., 2023; Pal et al., 2024).
Contribution. We first discover that the current practice of SNF initialization is highly suboptimal in terms of the training speed (Sitzmann et al., 2020b). Precisely, we find that one can accelerate the training speed of SNFs by up to 10$\times$, simply by scaling up the initial weights of all layers by a well-chosen factor. This extremely simple tuning, coined weight scaling (WS), provides a much greater speedup than the conventional SNF tuning strategy, i.e., scaling up the frequency multiplier of the activation functions (called $\omega_{0}$ in Sitzmann et al. (2020b)). Furthermore, this speedup can be achieved without any degradation in the generalization capabilities of the model, if we keep the scaling factor at a moderate level. Interestingly, experiments suggest that the optimal scaling factor, which strikes the balance between speedup and generalization, mainly depends on the physical scale of the NF workload (e.g., data resolution and model size) and remains similar over different data.
@@
Figure 1: A simple weight scaling accelerates training. The proposed weight scaling scales up the initial weights of an SNF by the factor of $\alpha$, except for the last layer (left panel). This scaling significantly reduces the number of steps required to achieve the desired level of training loss (right panel).
@@
To demystify why the weight scaling accelerates the SNF training, we conduct an in-depth analysis. In particular, we focus on understanding how the effects of WS on SNFs are similar yet different from ReLU nets, where the so-called lazy training phenomenon takes place (Chizat et al., 2019)—weight-scaled ReLU nets behave similarly to kernels, enjoying exponential convergence in theory for shallow nets but failing to achieve good accuracy at practical scales due to ill-conditioned trajectories. Our theoretical and empirical analyses highlight several key differences that sinusoidal activations bring, such as: (1) Weight-scaled SNFs also preserve activation distributions over layers, facilitating the gradient propagation when training a deep model (3.1). (2) Weight scaling increases not only the frequency of each basis, but also increases the relative magnitude of higher-order harmonics, which helps fit the high-frequency components abundant in natural signals (Lemmas 4.2 and 4.2). (3) Weight-scaled SNFs enjoy even better-conditioned trajectories than the unscaled models, as captured by the eigenspectrum analysis (Section 4.3).
@@ 2 Related work
Training efficiency of neural fields. There are two prevailing strategies to accelerate NF training. The first approach is to adopt advanced NF architectures, which mitigate the spectral bias of neural nets that hinders fitting high-frequency signals (Rahaman et al., 2019; Xu et al., 2020). Popular examples include utilizing specially designed positional encodings (Tancik et al., 2020; Müller et al., 2022) or activation functions (Sitzmann et al., 2020b; Ramasinghe & Lucey, 2022; Liu et al., 2024). Another strategy is to amortize the training cost via meta-learning, e.g., by training a fast-trainable initialization (Sitzmann et al., 2020a; Tancik et al., 2021) or constructing an auxiliary model that predicts NF parameters from the datum (Chen & Wang, 2022). This work aims to accelerate NF training through a better initialization. Unlike meta-learned initializations, however, our aim is to understand how the initialization (its scale, specifically) can facilitate training by analyzing how it affects optimization dynamics, without meta-learning involved.
Neural network initialization. Early works on network initialization focus on preserving the signals through layers to facilitate training deep networks. This is typically done via principled scaling of the initial weights (Glorot & Bengio, 2010; He et al., 2015), or by ensuring isometries in each layer through orthogonalization (Xiao et al., 2018). Recent works focus on the relationship between the initialization and the implicit bias of SGD (Vardi, 2023). In particular, the initialization scale determines whether the training takes place in the rich regime or kernel regime, which correspond to stronger and weaker influence of implicit bias, respectively (Woodworth et al., 2020; Varre et al., 2023). Empirically, weak implicit bias leads to the models trained from larger initialization to generalize poorly (Mehta et al., 2021). Our analyses share some spirit with the latter line of the literature. Distinct from the works, however, we put our main focus on how initialization affects the training speed, specifically targeting the neural fields.
Initializing sinusoidal neural fields. There is a limited number of works that study the initialization for SNFs. Sitzmann et al. (2020b) proposes an initialization scheme for SNF, designed in the signal propagation perspective. Tancik et al. (2021) develops a meta-learning algorithm to learn initializations for SNFs. Lee et al. (2021) extends the algorithm for sparse networks. Most related to our work, Saratchandran et al. (2024) proposes an initialization scheme that aims to facilitate the parameter-efficiency of SNFs. Different from this work, our work primarily focuses on the training efficiency, and achieves much faster training speed empirically.
3 Neural fields initialization via weight scaling
@@ 3.1 Network setup and notations
The SNF is trained using the gradient descent, with the mean-squared error (MSE) as a loss function. We assume that the input coordinates of the training data (i.e., the column vectors of the input data) have been scaled to lie inside the hypercube $[-1,+1]^{d_{0}}$, as in Sitzmann et al. (2020b).
@@ 3.3 Initialization with weight scaling
where we let $\alpha=1$ for the last layer; this design is due to the fact that any scaling in the last layer weight leads to a direct scaling of the initial function $f\mapsto\alpha f$, which is known to be harmful for the generalizability, according to the theoretical work of Chizat et al. (2019). Empirically, we observe that this weight scaling is much more effective in speeding up the training than a recently proposed variant which scales up only the last layer of the NF (Saratchandran et al., 2024); see Section 5.
@@
Proposition 3.1 (Informal; extension of Sitzmann et al. (2020b, Theorem 1.8)).
@@
The proposition implies that we can tune the scaling factor $\alpha$ within the range $[1,\infty)$ without any concerns regarding preserving the forward propagation signal. Thus, we can safely adopt the weight scaling for training a very deep sinusoidal neural field.
@@
The Fig. 2 illustrates this phenomenon for the case of image regression, where we train an SNF to fit a downsampled image until some desired peak signal-to-noise ratio (PSNR) is met, and then use the model to interpolate between seen pixel coordinates.
We observe that by increasing $\alpha$ from $1$, the loss on unseen coordinates increase, and the training speed increases as well. Notably, there exists some $\alpha>1$ where the test PSNR remains similar while the training speed has strictly increased.
@@ 4 Understanding the effects of the weight scaling
Comparison with lazy training. At the first glance, it is tempting to argue that this phenomenon is connected to the “lazy training” phenomenon (Chizat et al., 2019; Woodworth et al., 2020): Scaling the initial weights ($\mathbf{W}\mapsto\alpha\mathbf{W}$) amplifies the initial model functional ($f\mapsto\alpha^{l}f$), which in turn leads to the model to be linearized; then, one can show that the SGD converges at an exponential rate, i.e., the fast training, with the trained parameters being very close to the initial one. However, there is a significant gap between this theoretical understanding and our observations for the SNFs:
  • The proposed weight scaling does not amplify the initial functional, as sinusoidal activations are not positively homogeneous (i.e., $\sigma(ax)\neq a\cdot\sigma(x)$, even for positive $a,x$), unlike ReLU. In fact, WS keeps the last layer parameters unscaled, making it further from any functional amplification, and empirically works better than a direct amplification of functional (Saratchandran et al., 2024).
@@ 4.2 Frequency spectrum of the initial functional
How about for deeper networks? Our theoretical and empirical analyses reveals that, for deeper SNFs, the weight scaling also increases the relative power of the higher-frequency bases. Thus, the weight-scaled networks may suffer less from the spectral bias, and can easily represent signals with large high-frequency components (Rahaman et al., 2019).
@@
Lemma 4.2 (Scaling of harmonics).
@@
Empirical analysis. To demonstrate that the increase of coefficients in the higher-order harmonics takes place indeed, we conduct a spectral analysis of an initialized SNF for audio data (Fig. 4). From the plot, we observe that the initialized SNFs with larger scaling factor $\alpha$ has a richer higher frequency spectrum. Moreover, the decay rate of the higher-frequency components, as we consider increase the frequency level, is much smaller for larger $\alpha$; as a consequence, weight-scaled SNFs tend to have non-zero coefficients for very high frequency ranges, enabling a faster fitting to the natural audio signals (gray) with rich high-frequency components.
@@ 4.3 Optimization trajectories of weight-scaled networks
Condition number. Roughly, the condition number denotes the ratio $\lambda_{\max}/\lambda_{\min}$ of the largest and the smallest eigenvalues of the eNTK; this quantity determines the exponent of the exponential convergence rates of the kernelized models, with larger condition number meaning the slower training (Chizat et al., 2019). In Fig. 5(a), we plot how the condition numbers of the eNTK evolve during the training for SNFs, under weight scaling or frequency tuning; for numerical stability, we take an average of top-5 and bottom-5 eigenvalues and take their ratio. We observe that the weight-scaled SNFs tend to have a much smaller condition number at initialization, which becomes even smaller during the training. On the other hand, the SNF initialized with the standard initialization scheme suffers from a very high condition number, which gets even higher during the training.
@@
where $\phi_{i}$ denotes the $i$-th eigenvector of the kernel and $\mathbf{e}=Y-f(X)$ denotes the vector of prediction errors in the pixels (Kopitkov & Indelman, 2020). This quantity measures how much of the residual signal is concentrated in the directions of large eigenvalues, which is theoretically easier to be expressed or optimized by kernel learning (Baratin et al., 2021; Yüce et al., 2022). In Fig. 5(b), we compare the kernel-task alignments during the SNF training trajectories. We observe that the weight scaling gives a highly optimizable kernel throughout all stages of training.
@@ 5.1 Main experiments
Xavier Uniform (Glorot & Bengio, 2010) Sinusoidal 0.46$\pm$0.10 0.39$\pm$0.10 0.0000$\pm$0.0000 4.11$\pm$0.66 7.77$\pm$0.20
ReLU + P.E. (Mildenhall et al., 2020) ReLU 18.60$\pm$0.08 16.72$\pm$0.08 0.9896$\pm$0.0003 33.30$\pm$0.54 24.98$\pm$0.19
FFN (Tancik et al., 2020) ReLU 20.52$\pm$0.60 19.81$\pm$0.48 0.9843$\pm$0.0020 38.69$\pm$0.27 16.66$\pm$0.28
SIREN init. (Sitzmann et al., 2020b) Sinusoidal 24.58$\pm$0.05 22.86$\pm$0.06 0.9925$\pm$0.0001 38.72$\pm$0.07 37.37$\pm$3.11
GaussNet (Ramasinghe & Lucey, 2022) Gaussian 21.94$\pm$2.48 19.22$\pm$0.14 0.9914$\pm$0.0005 38.56$\pm$0.51 27.47$\pm$2.10
WIRE (Saragadam et al., 2023) Wavelet 28.94$\pm$0.21 28.20$\pm$0.13 0.9912$\pm$0.0005 31.27$\pm$0.53 16.83$\pm$1.85
NFSL (Saratchandran et al., 2024) Sinusoidal 24.93$\pm$0.07 23.39$\pm$0.09 0.9925$\pm$0.0001 38.92$\pm$0.07 37.17$\pm$2.88
Weight scaling (ours) Sinusoidal 42.83$\pm$0.35 42.03$\pm$0.41 0.9941$\pm$0.0002 45.28$\pm$0.03 45.04$\pm$3.23
We validate the effectiveness of WS in different data domains by comparing against various neural fields. To compare the training speed, we fix the number of training steps and compare the training accuracy. In particular, we consider the following tasks and baselines.
Task: Image regression. The network is trained to approximate the signal intensity $c$, for each given normalized 2D pixel coordinates $(x,y)$. For our experiments, we use Kodak (Kodak, 1999) and DIV2K (Agustsson & Timofte, 2017) datasets. Each image is resized to a resolution of $512\times 512$ in grayscale, following Lindell et al. (2022); Seo et al. (2024). We report the training PSNR after a full-batch training for 150 iterations. For all NFs, we use five layers with width 512.
@@
Baselines. We compare the weight scaling against seven baselines; we have selected the memory-efficient and versatile neural fields, rather than highly specialized and heavy ones, such as Instant-NGP (Müller et al., 2022). More concretely, we use the following baselines: (1) Xavier uniform: the method by (Glorot & Bengio, 2010) applied on SNF, (2) SIREN init.: SNF using the standard initialization of SIREN (Sitzmann et al., 2020b), (3) NFSL: A recent initialization scheme motivated by scaling law (Saratchandran et al., 2024). (4) ReLU+P.E.: ReLU nets with positional encodings (Mildenhall et al., 2020). (5) FFN: Fourier feature networks Tancik et al. (2020). (6) GaussNet: MLP with Gaussian activations (Ramasinghe & Lucey, 2022). (7) WIRE: MLP with Gabor wavelet-based activations (Saragadam et al., 2023). The baselines (1-3) use SNFs as the backbone architecture, while (4-7) uses other architectures with different activation functions.
Other details. For selecting the hyperparameters for each NFs (including $\alpha$), we have conducted a grid search. We use the Adam optimizer (Kingma, 2015) with a fixed learning rate specific to each data domain. Other experimental details are provided in Appendix E.
Result. Table 1 reports the comprehensive results across various data domains. We observe that the weight scaling significantly outperforms all baselines throughout all tasks considered. In Section E.4, we provide further qualitative results on various modalities, where we find that the weight scaling enhances capturing the high frequency details. These results support our hypothesis that the richer frequency spectrum provided by the weight scaling mitigates the spectral bias (Section 4.2), thereby enabling a faster training of neural fields.
Additional experiments. We have also conducted experiments on the NF-based dataset construction tasks (Papa et al., 2024). We report the results in Section E.2.

### v3

@@ Abstract
[…]
@@ 1 Introduction
For neural fields, the training speed is of vital importance. Representing each datum as an NF requires tedious training of a neural network, which can take up to several hours of GPU-based training (Mildenhall et al., 2020). This computational burden becomes a critical obstacle toward adoption to tasks that involve large-scale data, such as NF-based inference or generation (Ma et al., 2024; Papa et al., 2024). To address this issue, many research have been taken to accelerate NF training, including fast-trainable NF architectures (Sitzmann et al., 2020b; Müller et al., 2022), meta-learning (Sitzmann et al., 2020a; Chen and Wang, 2022), and data transformations (Seo et al., 2024).
@@
Despite such efforts, a critical aspect of NF training remains understudied: How should we initialize NFs for the fastest training? While the importance of a good initialization has been much contemplated in classic deep learning literature (Glorot and Bengio, 2010; Zhang et al., 2019), such understandings do not immediately answer our question, for many reasons. First, conventional works mostly focus on how initialization affects the signal propagation properties of deep networks with hundreds of layers, while most NFs use shallow models with only a few layers (Mildenhall et al., 2020). Second, conventional works aim to maximize the model quality at convergence, while in NFs it is often more important how fast the network achieves certain fidelity criterion (Müller et al., 2022). Lastly, NFs put great emphasis on how the model performs on seen inputs (i.e., training loss) and leaves how well it interpolates between coordinates (i.e., generalizability) as a secondary issue, while conventional neural networks care exclusively on the test performance.
@@
To fill this gap, in this paper, we investigate how the initialization affects the training efficiency of neural fields. In particular, we focus on initializing sinusoidal neural field (SNF), i.e., multilayer perceptrons with sinusoidal activation functions (Sitzmann et al., 2020b). SNF is one of the most widely used NFs, due to its versatility and parameter-efficiency; the architecture works as a strong baseline in a wide range of data modalities (Grattarola and Vandergheynst, 2022; Kim et al., 2024; Rußwurm et al., 2024) and is used as an indispensable building block of many state-of-the-art neural field frameworks (Guo et al., 2023; Schwarz et al., 2023; Pal et al., 2024).
[…]
@@
Figure 1: A simple weight scaling accelerates training. The proposed weight scaling scales up the initial weights of an SNF by the factor of $\alpha$, except for the last layer (left panel). The weight scaled SNF significantly speeds up training across a variety of methods (right panel: train PSNR curve for a single Kodak image).
@@
To demystify why the weight scaling accelerates the SNF training, we conduct an in-depth analysis. In particular, we focus on understanding how the effects of WS on SNFs are similar yet different from ReLU nets, where the so-called lazy training phenomenon takes place (Chizat et al., 2019)—weight-scaled ReLU nets behave similarly to kernels, enjoying exponential convergence in theory for shallow nets but failing to achieve good accuracy at practical scales due to ill-conditioned trajectories. Our theoretical and empirical analyses highlight several key differences that sinusoidal activations bring, such as: (1) Weight-scaled SNFs also preserve activation distributions over layers, facilitating the gradient propagation when training a deep model (1). (2) Weight scaling increases not only the frequency of each basis, but also increases the relative magnitude of higher-order harmonics, which helps fit the high-frequency components abundant in natural signals (Lemmas 3 and 4.2). (3) Weight-scaled SNFs enjoy even better-conditioned trajectories than the unscaled models, as captured by the eigenspectrum analysis (Section 4.3).
@@ 2 Related work
Training efficiency of neural fields. There are two prevailing strategies to accelerate NF training. The first approach is to adopt advanced NF architectures (Fathony et al., 2021; Dou et al., 2023), which mitigate the spectral bias of neural nets that hinders fitting high-frequency signals (Rahaman et al., 2019; Basri et al., 2019; Xu et al., 2020; Cao et al., 2021; Choraria et al., 2022). Popular examples include utilizing specially designed positional encodings (Tancik et al., 2020; Müller et al., 2022) or activation functions (Sitzmann et al., 2020b; Ramasinghe and Lucey, 2022; Liu et al., 2024; Saratchandran et al., 2024c). Another strategy is to amortize the training cost via meta-learning, e.g., by training a fast-trainable initialization (Sitzmann et al., 2020a; Tancik et al., 2021) or constructing an auxiliary model that predicts NF parameters from the datum (Chen and Wang, 2022). This work aims to accelerate NF training through a better initialization. Unlike meta-learned initializations, however, our aim is to understand how the initialization (its scale, specifically) can facilitate training by analyzing how it affects optimization dynamics, without meta-learning involved.
Neural network initialization. Early works on network initialization focus on preserving signals through layers via principled weight scaling (Glorot and Bengio, 2010; He et al., 2015) or orthogonalization (Xiao et al., 2018). Recent studies explore its impact on the implicit bias of SGD (Vardi, 2023), where initialization scale determines whether training occurs in the rich or kernel regime (Jacot et al., 2018; Woodworth et al., 2020; Varre et al., 2023). While aligned with this line of research, we primarily investigate how initialization affects training speed, particularly in neural fields.
Initializing sinusoidal neural fields. There is a limited number of works that study the initialization for SNFs. Sitzmann et al. (2020b) proposes an initialization scheme for SNF, designed in the signal propagation perspective. Tancik et al. (2021) develops a meta-learning algorithm to learn initializations for SNFs. Lee et al. (2021) extends the algorithm for sparse networks. Most related to our work, Saratchandran et al. (2024b) proposes an initialization scheme that aims to facilitate the parameter-efficiency of SNFs. Different from this work, our work primarily focuses on the training efficiency, and achieves much faster training speed empirically.
Sinusoidal activation functions in ML. Beyond their well-known use in NF literature (Sitzmann et al., 2020b; Ashkenazi and Treister, 2024), sinusoidal activation functions have also been studied in other domains. Examples include Fourier neural networks (Gashler and Ashmore, 2014), Bayesian neural networks and out-of-distribution detection (Meronen et al., 2021), low-rank learning (Ji et al., 2025), neural compression (Thrash et al., 2025), and generative models (Chan et al., 2021).
3 Neural field initialization via weight scaling
@@ 3.1 Network setup and notations
The SNF is trained using the gradient descent, with the mean-squared error (MSE) as a loss function. We assume that the input coordinates of the training data (i.e., the column vectors of the input data) have been scaled to lie inside the hypercube $1,+1]^{d_{0}}$, as in Sitzmann et al. (2020b).
@@ 3.3 Initialization with weight scaling
where we let $\alpha=1$ for the last layer; this design is due to the fact that any scaling in the last layer weight leads to a direct scaling of the initial function $f\mapsto\alpha f$, which is known to be harmful for the generalizability, according to the theoretical work of Chizat et al. (2019). Empirically, we observe that this weight scaling is much more effective in speeding up the training than a recently proposed variant which scales up only the last layer of the NF (Saratchandran et al., 2024b); see Section 5.
@@
Proposition 1 (Informal; extension of Sitzmann et al. (2020b, Theorem 1.8)).
@@
The proposition implies that we can tune the scaling factor $\alpha$ within the range $1,\infty)$ without any concerns regarding preserving the forward propagated signal. Thus, we can safely adopt the weight scaling for training a very deep sinusoidal neural field.
@@
The Fig. 2 illustrates this phenomenon for the case of image regression, where we train an SNF to fit a downsampled image until some desired peak signal-to-noise ratio (PSNR) is met, and then use the model to interpolate between seen pixel coordinates. We observe that by increasing $\alpha$ from $1$, the loss on unseen coordinates increase, and the training speed increases as well. Notably, there exists some $\alpha>1$ where the test PSNR remains similar while the training speed has strictly increased.
@@ 4 Understanding the effects of the weight scaling
Comparison with lazy training. At the first glance, it is tempting to argue that this phenomenon is connected to the “lazy training” (Chizat et al., 2019; Woodworth et al., 2020): Scaling the initial weights ($\mathbf{W}\mapsto\alpha\mathbf{W}$) amplifies the initial model functional ($f\mapsto\alpha^{l}f$) (Taheri et al., 2021), which in turn leads to the model to be linearized; then, one can show that the SGD converges at an exponential rate, i.e., the fast training, with the trained parameters being very close to the initial one. However, there is a significant gap between this theoretical understanding and our observations for the SNFs:
  • The proposed weight scaling does not amplify the initial functional, as sinusoidal activations are not positively homogeneous (i.e., $\sigma(ax)\neq a\cdot\sigma(x)$, even for positive $a,x$), unlike ReLU. In fact, WS keeps the last layer parameters unscaled, making it further from any functional amplification, and empirically works better than a direct amplification of functional (Saratchandran et al., 2024b).
@@ 4.2 Frequency spectrum of the initial functional
How about for deeper networks? Our theoretical and empirical analyses reveal that, for deeper SNFs, the weight scaling also increases the relative power of the higher-frequency bases. Thus, the weight-scaled networks may suffer less from the spectral bias, and can easily represent signals with large high-frequency components (Rahaman et al., 2019).
@@
Lemma 3 (Scaling of harmonics).
@@
Empirical analysis. To demonstrate that the increase of coefficients in the higher-order harmonics takes place indeed, we conduct a spectral analysis of an initialized SNF for natural image (Fig. 4). From the plot, we observe that the initialized SNFs with larger scaling factor $\alpha$ has a richer higher frequency spectrum. Moreover, the decay rate of the higher-frequency components, as we consider increase the frequency level, is much smaller for larger $\alpha$; as a consequence, weight-scaled SNFs tend to have non-zero coefficients for very high frequency ranges, enabling a faster fitting to the natural audio signals (gray) with rich high-frequency components.
@@ 4.3 Optimization trajectories of weight-scaled networks
Condition number. Roughly, the condition number denotes the ratio $\lambda_{\max}/\lambda_{\min}$ of the largest and the smallest eigenvalues of the eNTK; this quantity determines the exponent of the exponential convergence rates of the kernelized models, with larger condition number meaning the slower training (Chizat et al., 2019). In Fig. 5(a), we plot how the condition numbers of the eNTK evolve during the training for SNFs, under weight scaling or frequency tuning; for numerical stability, we take an average of top-5 and bottom-5 eigenvalues and take their ratio. We observe that the weight-scaled SNFs tend to have a much smaller condition number at initialization, which becomes even smaller during training. On the other hand, the SNF with the standard initialization suffers from a high condition number, which gets even higher during training. More discussion can be found in Section C.3.
@@
where $\phi_{i}$ denotes the $i$-th eigenvector of the kernel and $\mathbf{e}=Y-f(X)$ denotes the vector of prediction errors in the pixels (Kopitkov and Indelman, 2020). This quantity measures how much of the residual signal is concentrated in the directions of large eigenvalues, which is theoretically easier to be expressed or optimized by kernel learning (Baratin et al., 2021; Yüce et al., 2022). In Fig. 5(b), we compare the kernel-task alignments during the SNF training trajectories. We observe that the weight scaling gives a highly optimizable kernel throughout all stages of training.
@@ 5.1 Main experiments
[…]
We validate the effectiveness of WS in different data domains by comparing against various neural fields. To compare the training speed, we compare the training accuracy for equivalent steps. In particular, we consider the following tasks and baselines. Other details can be found in Section F.4.
Task: Image regression. The network is trained to approximate the signal intensity $c$, for each given normalized 2D pixel coordinates $(x,y)$. For our experiments, we use Kodak (Kodak, 1999) and DIV2K (Agustsson and Timofte, 2017) datasets. Each image is resized to a resolution of $512\times 512$ in grayscale, following Lindell et al. (2022); Seo et al. (2024). We report the training PSNR after a full-batch training for 150 iterations. For all NFs, we use five layers with width 512.
@@
Baselines. We compare the weight scaling against seven baselines; we have selected the versatile neural field methods, rather than highly specialized and heavy ones, such as Instant-NGP (Müller et al., 2022). More concretely, we use the following baselines: (1) Xavier uniform: the method by (Glorot and Bengio, 2010) applied on SNF, (2) SIREN init.: SNF using the standard initialization of SIREN (Sitzmann et al., 2020b), (3) NFSL: A recent initialization scheme motivated by scaling law (Saratchandran et al., 2024b). (4) ReLU+P.E.: ReLU nets with positional encodings (Mildenhall et al., 2020). (5) FFN: Fourier feature networks (Tancik et al., 2020). (6) MFN: multiplicative filter networks (Fathony et al., 2021). (7) GaussNet: MLP with Gaussian activations (Ramasinghe and Lucey, 2022). (8) WIRE: MLP with Gabor wavelet-based activations (Saragadam et al., 2023). (9) Sinc NF: MLP with sinc (i.e., $\sin(\omega x)/\omega x$) activations (Saratchandran et al., 2024c). The baselines (1-3) use SNFs, while (4-9) uses other architectures with different activation functions.
Result. Table 1 reports the comprehensive results across various data domains. We observe that the weight scaling significantly outperforms all baselines throughout all tasks considered. In Section F.8, we provide further qualitative results on various modalities, where we find that the weight scaling enhances capturing the high frequency details. These results support our hypothesis that the richer frequency spectrum provided by the weight scaling mitigates the spectral bias (Section 4.2), thereby enabling a faster training of neural fields.
Additional experiments. We have also conducted experiments on the NF-based dataset construction tasks (Papa et al., 2024), novel view synthesis via neural radiance fields (NeRFs) (Mildenhall et al., 2020), and solving partial differential equations. We report the results and details for each experiment in Section F.5, Section F.6, and Section F.7 respectively.